News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Wednesday 27 October 1999 10.00am EST

Yesterday the Chandra team in Cambridge was visited by Commander Eileen Collins and her crew together with Fred Wojtalik and Martin Weisskopf from MSFC. The crew showed an extraordinary HDTV movie of the launch and deploy, made presentations and spent time talking with the team.

During the last 48 hours, Chandra completed GTO observations of E0102-72.3, NGC2516, and Sirius-B, performed a series of calibration observations of LMC X-1 with the HRC, and passed through the radiation belts. During the next 24 hours observations will be conducted of Sirius-B with HRC-S/LETG, and HR1099 and 2REJ1032+532 with HRC-I.

Quick-look analysis of HRC data from the last 24 hours indicated several observations with HRC-S had been completed without fiducial light acquisition. The problem was traced to a ground system software error when transitioning between HRC-I and HRC-S observations with the same numbered set of fid-lights. The problem has been corrected in the next load, and a ground system software patch put in work.

A flight software patch to modify the HRC voltage step values was uplinked yesterday. The voltage step values are based on on-orbit data, and will optimize the HRC imaging performance.

A real-time procedure to collect ACIS CCD data in raw mode was completed overnight. A second test is planned for next week. These tests are related to CCD testing performed at GSFC last week as part of the effort to understand the cause of the CTI degradation.

The spacecraft systems continued to operate nominally.





Return to Status Reports